OpenStudy (anonymous):

Holly missed 7 of 181 school days this year. Approximately what percent of school days was Holly in attendance? a 26% b 74% c 96% d 4%

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

@itsthatguy

OpenStudy (anonymous):

@itsthatguy

OpenStudy (anonymous):

@Ilovecake

OpenStudy (anonymous):

@globule

OpenStudy (anonymous):

@newwar248

OpenStudy (anonymous):

@gee

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@undeadknight26

geerky42 (geerky42):

She missed 7 days so that means she didn't miss 181 - 7 =174 days of school. So you have \(\dfrac{\text{Days attended }}{\text{Total days}}\times100 ~~=~~ \dfrac{174}{181}\times100 = \underline{~~~?~~~}\%\)